# GenoTEX: A Benchmark for Automated Gene Expression Data Analysis in Alignment with Bioinformaticians

### v2

@@ 6 Conclusion
In this paper, we introduce GenoTEX, a benchmark for automated gene expression data analysis aimed at identifying disease-associated genes. The benchmark represents the standard pipeline in computational genomics, with problems and solutions curated by expert bioinformaticians. By structuring evaluation around three tasks—dataset selection, data preprocessing, and statistical analysis—GenoTEX provides a framework for assessing automated computational approaches to gene identification.
@@
  • Appendix D outlines the criteria for forming trait-condition pairs for gene identification problems in our benchmark.
@@
  • Appendix F describes our statistical method for gene identification.
@@ A.1 Documentation and intended uses
GenoTEX is documented following the Datasheets for Datasets [26] framework, providing a comprehensive description of the data collection process, preprocessing steps, and statistical analysis methods employed. The detailed datasheet is available here. Each step of the pipeline aims to mirror the standards of computational genomics, ensuring that the dataset is both accurate and reliable. The intended uses of GenoTEX are broad, encompassing the evaluation and development of AI-driven methods for genomics data analysis. By providing a standardized benchmark, GenoTEX aims to facilitate the advancement of machine learning models capable of automating the complex task of gene expression analysis. Researchers in bioinformatics and related fields can leverage this dataset to benchmark their algorithms, fostering innovation and improving the scalability of gene identification processes.
@@ A.4 Data format and persistent identifiers
GenoTEX is provided in open and widely used data formats, including CSV and JSON, ensuring compatibility with a wide range of analytical tools and platforms. Detailed instructions on how to read and use the dataset are included in the documentation, making it accessible to both novice and experienced researchers. To enhance the dataset’s stability and ease of reference, we have minted a DOI for GenoTEX. This DOI will serve as a reliable means of access and citation for the dataset, promoting its use in academic and professional research.
@@ Appendix B Guidelines for gene expression data analysis
c. Choose the appropriate data type (continuous, binary, or categorical) and design conversion functions to encode the features into that type.
@@
1. Select the best input data relevant to the gene identification problem, and load the data into a dataframe. If multiple preprocessed datasets are available for statistical analysis about a trait, we select the one with the largest sample size.
@@
1. Determine the variable type (binary, continuous, or categorical) of the condition.
@@ Appendix C Examples of manual analysis
In addition to the guidelines file, we provide example files to the participants of our benchmark curation. These examples include code and results for analyzing gene identification problems related to traits such as Breast Cancer and Epilepsy. These illustrations have proven helpful in familiarizing participants with these tasks quickly. Among the many steps in the analysis pipeline, a key step is the trait data extraction during the preprocessing of GEO data. This step requires biomedical knowledge and an understanding of the dataset collection process described in the metadata. In this section, we will introduce the part of the manual analysis examples related to this crucial step.
@@ Appendix D Criteria for manual correction of trait-condition pairs
To ensure the scientific validity of our benchmark problems, we apply specific rules for including and excluding certain trait-condition pairs. Each biomedical entity in our list can be considered a trait and paired with a condition, where the condition is either another entity from the list or a demographic attribute like "age" or "gender." The following criteria are designed to maintain scientific relevance and robustness:
  • Trait-Condition Role Assignment: Entities such as language abilities, Vitamin D levels, and bone density are included only as conditions and not as traits. This distinction ensures that the primary focus remains on traits with more direct clinical implications, while these entities serve as influential factors that could affect those traits.
  • Universal Conditions: Entities such as obesity, hypertension, and mental disorders like anxiety disorder and bipolar disorder are designated as conditions to be paired with all other traits. This is because these conditions are widespread and significantly impact various health outcomes, making them critical factors to consider in any genetic analysis.
  • Gender-Specific Considerations: Gender-specific entities such as prostate cancer, endometriosis, and breast cancer are not conditioned on gender. Furthermore, entities from different genders are not paired. This approach respects the biological distinctions between genders and ensures that the resulting problems remain relevant and meaningful.
@@
These criteria are used in combination with the Jaccard similarity of related genes (Section 3.2), to uphold the scientific integrity and relevance of the benchmark problems, facilitating meaningful and insightful gene expression analysis.
Appendix E Details about the data sources
GEO
The Gene Expression Omnibus (GEO) [14] is a public archive for high-throughput gene expression data and various other types of genomic data. We leveraged the Entrez programming utility to perform a systematic search of the GEO database for human series data relevant to each trait on our list, prioritizing datasets with large sample sizes. We downloaded both SOFT and matrix files for each series and used heuristic evaluations of file sizes to pinpoint datasets likely containing gene expression data. When automated searches failed to yield results for specific traits, we conducted manual searches using expanded synonyms from Medical Subject Headings (MeSH) terms.
TCGA-Xena
The Cancer Genome Atlas (TCGA) [72], accessed through the Xena platform [28], offers a rich repository of RNAseq gene expression and clinical data for numerous cancer types. We obtained data for 36 traits from the TCGA cohort using the UCSC Xena platform, which provides high-quality, cancer-related gene expression and clinical data linked by patient IDs.
NCBI Gene
The NCBI Gene database [10] is an important resource for comprehensive information on gene sequences, functions, and their links to diseases and conditions. For each trait, we queried the database to compile sets of gene symbols associated with the trait. This data was crucial for identifying disease-disease associations for problem generation and for selecting common regressors in two-step regression analyses.
Appendix F Details of statistical method for gene identification
@@
After the detection of confounding factors, if the agent believes there are no confounding factors in the data, it will directly follow the vanilla case above. However, if there are confounding factors, the agent will need to decide one of the following two approaches.
@@
Note that linear mixed model will estimate the coefficients individually for each gene, which is not necessarily ideal for our analysis. As an alternative choice, Statistician can also use the following.
Regression after Confounding Factor Correction In order to have a multivariate method, where the coefficients of genes can be considered together, the Statistician can also choose to correct the confounding factors first. The correction is mostly built upon the theory of how linear mixed model helps to correct confounding factors [76], and achieved by the rotation of $\mathbf{X}$ as following:
@@ Appendix G Challenges faced by existing methods on our benchmark
We tested these methods under three different settings: (i) providing general task instructions, (ii) providing detailed task instructions used by GenoAgent, and (iii) providing detailed task instructions and all necessary library functions as in GenoAgent. Each setting was tested on a subset of 50 gene identification problems. Our results show that none of the methods generated runnable code for preprocessing datasets downloaded from GEO. Persistent errors in the generated code prevented testable outputs, regardless of the level of detail provided.
@@ Appendix H Discussion on the limitations of GenoAgent
This section discusses the observed limitations of our baseline method, GenoAgent, on the GenoTEX benchmark. We identified that certain steps are inherently challenging, and instability in the feedback mechanism may hinder the agents’ iterative improvement process. Figures 5 and 5 illustrate the two types of feedback mechanisms in GenoAgent.

### v3

@@ 6 Conclusion
In this paper, we introduce GenoTEX, a benchmark for automated gene expression data analysis aimed at identifying disease-associated genes. The benchmark represents the standard pipeline in computational genomics, with problems and solutions curated by expert bioinformaticians. By structuring evaluation around three tasks—dataset selection, data preprocessing, and statistical analysis—GenoTEX provides a framework for assessing automated computational approaches to GTA analysis.
@@
  • Appendix D outlines the criteria for forming trait-condition pairs for GTA analysis problems in our benchmark.
@@
  • Appendix F describes our statistical method for GTA analysis.
@@ A.1 Documentation and intended uses
GenoTEX is documented following the Datasheets for Datasets [26] framework, providing a comprehensive description of the data collection process, preprocessing steps, and statistical analysis methods employed. The detailed datasheet is available here. Each step of the pipeline aims to mirror the standards of computational genomics, ensuring that the dataset is both accurate and reliable. The intended uses of GenoTEX are broad, encompassing the evaluation and development of AI-driven methods for genomics data analysis. By providing a standardized benchmark, GenoTEX aims to facilitate the advancement of machine learning models capable of automating the complex task of gene expression analysis. Researchers in bioinformatics and related fields can leverage this dataset to benchmark their algorithms, fostering innovation and improving the scalability of gene-trait association (GTA) analysis processes.
@@ A.4 Data format and persistent identifiers
GenoTEX is provided in open and widely used data formats, including CSV and JSON, ensuring compatibility with a wide range of analytical tools and platforms. Detailed instructions on how to read and use the dataset are included in the documentation, making it accessible to both novice and experienced researchers. To enhance the dataset’s stability and ease of reference, we have created a DOI for GenoTEX: https://doi.org/10.34740/kaggle/dsv/11309048. This DOI will serve as a reliable means of access and citation for the dataset, promoting its use in academic and professional research.
@@ Appendix B Guidelines for gene expression data analysis
c. Choose the appropriate data type (continuous or binary) and design conversion functions to encode the features into that type.
@@
1. Select the best input data relevant to the GTA analysis problem, and load the data into a dataframe. If multiple preprocessed datasets are available for statistical analysis about a trait, we select the one with the largest sample size.
@@
1. Determine the variable type (continuous or binary) of the condition.
@@ Appendix C Examples of manual analysis
In addition to the guidelines file, we provide example files to the participants of our benchmark curation. These examples include code and results for analyzing GTA analysis problems related to traits such as Breast Cancer and Epilepsy. These illustrations have proven helpful in familiarizing participants with these tasks quickly. Among the many steps in the analysis pipeline, a key step is the trait data extraction during the preprocessing of GEO data. This step requires biomedical knowledge and an understanding of the dataset collection process described in the metadata. In this section, we will introduce the part of the manual analysis examples related to this crucial step.
@@ Appendix D Criteria for manual correction of trait-condition pairs
To ensure the scientific validity of our benchmark problems, we apply specific rules for including and excluding certain trait-condition pairs. Each biomedical entity in our list can be considered a trait and paired with a condition, where the condition is either another entity from the list or a demographic attribute like “age” or “gender”. The following criteria are designed to maintain scientific relevance and robustness:
  • Trait-Condition Role Assignment: Entities such as Vitamin D levels and LDL cholesterol levels are included only as conditions and not as traits. This distinction ensures that the primary focus remains on traits with more direct clinical implications, while these entities serve as influential factors that could affect those traits.
  • Universal Conditions: Entities such as obesity and hypertension are designated as conditions to be paired with all other traits. This is because these conditions are widespread and significantly impact various health outcomes, making them critical factors to consider in any genetic analysis.
  • Gender-Specific Considerations: Gender-specific entities such as endometriosis and prostate cancer are not conditioned on gender. Furthermore, entities from different genders are not paired. This approach respects the biological distinctions between genders and ensures that the resulting problems remain relevant and meaningful.
@@
These criteria are used in combination with the Jaccard similarity of associated genes (Section 3.2), to ensure the biological significance of the benchmark problems.
Appendix E Details about data acquisition
Our benchmark required two categories of data sources: (1) gene expression and clinical datasets as primary input data, and (2) domain knowledge repositories to support analysis and evaluation. Below, we detail our systematic approach to acquiring and processing data from each source.
E.1 Input Gene Expression Data Sources
Gene Expression Omnibus (GEO)
GEO [14] is a comprehensive public repository for high-throughput genomic data. We implemented a systematic acquisition protocol using the Entrez Programming Utilities to search GEO DataSets (a manually curated subset of GEO) for human gene expression series relevant to each trait in our study. Our selection process followed these criteria:
  1. 1.
    Search parameters:
    • Sample size: Between 30 and 10,000 subjects
    • Organism: Homo sapiens (human)
    • Publication period: 2010-2025
    • Data types: Expression profiling (array and high-throughput sequencing), genome variation profiling (high-throughput sequencing and SNP array), SNP genotyping, and third-party reanalysis
  2. 2.
    File requirements:
    • Presence of both SOFT (Series and Platform family) files and matrix files
    • Matrix file size: 0.1MB-100MB (ensuring sufficient gene expression data while maintaining computational tractability)
    • Family file size: Less than 100MB
  3. 3.
    Cohort selection:
    • For each trait, we selected up to 10 cohort datasets that satisfied all criteria
    • When more than 10 datasets were available, we used GEO’s default ranking system (which prioritizes recency and relevance)
This approach ensured a balanced representation of diverse, high-quality datasets of manageable size for each trait.
The Cancer Genome Atlas (TCGA)
The TCGA [76] provides standardized gene expression and clinical data for multiple cancer types. We accessed these datasets through the UCSC Xena platform [28], which offers harmonized data in a consistent format. Unlike our selective approach with GEO, we comprehensively incorporated all available TCGA datasets related to 36 cancer types for several reasons:
  1. 1.
    Standardized format: TCGA data through Xena has consistent processing, with uniformly formatted gene expression matrices and clinical information
  2. 2.
    High quality: The datasets undergo rigorous quality control through both TCGA and the Xena platform
  3. 3.
    Integrated patient IDs: Expression and clinical data are consistently linked through standardized patient identifiers
  4. 4.
    Appropriate size: Dataset sizes are consistently suitable for computational analysis without additional filtering
This comprehensive inclusion of TCGA data complemented our more selective approach to GEO datasets, providing our benchmark with standardized cancer genomics data.
E.2 Domain Knowledge Sources
NCBI Gene Database
To address nomenclature inconsistencies across expression datasets, we utilized the NCBI Gene database [10] to standardize gene symbols. Our process involved:
  1. 1.
    Extracting official human gene symbols and their synonyms from the NCBI Gene FTP repository
  2. 2.
    Creating a gene synonym dictionary mapping all gene aliases (converted to uppercase) to their respective official symbols
  3. 3.
    Implementing precedence rules to resolve ambiguities where multiple official symbols share the same synonym
This standardization facilitated the integration of expression data from heterogeneous sources and ensured consistent gene identification across all datasets in our benchmark.
Open Targets Platform
The Open Targets Platform [74] integrates multi-omics data to identify and prioritize disease-associated targets. We specifically utilized the “associationByOverallDirect” dataset, which contains direct target-disease associations with overall scores aggregated across all evidence types. Our extraction process included:
  1. 1.
    Downloading the association dataset from Open Targets using their rsync service
  2. 2.
    Mapping each trait name to a disease ID using the Open Targets API
  3. 3.
    Filtering associations to include only those with confidence scores exceeding 0.2 to ensure biological relevance
  4. 4.
    Converting target Ensembl IDs to gene symbols using mygene.info [90] and normalizing them using our gene synonym dictionary
These curated gene-disease associations served two crucial functions in our benchmark construction:
  1. 1.
    Selection of meaningful trait-condition pairs: We calculated Jaccard similarity between trait-related and condition-related gene sets, selecting pairs with similarity exceeding 0.1 for their shared biological mechanisms
  2. 2.
    Identification of biologically-relevant regressors: We identified genes common between trait and condition datasets that also had known associations with the condition according to Open Targets. This approach enabled us to train regression models on condition datasets and apply them to predict condition status in trait datasets, facilitating systematic study of how specific conditions influence gene-trait relationships. Section F provides more details on this “two-step regression” approach.
Appendix F Details of statistical method for GTA analysis
@@
After the confounding factor detection, if we observe no confounding factors in the data, we directly follow the vanilla case above. However, if there are confounding factors, we adopt one of the following two approaches.
@@
Note that linear mixed model will estimate the coefficients individually for each gene, which is not necessarily ideal for our analysis. As an alternative choice, we can also use the following.
Regression after Confounding Factor Correction In order to have a multivariate method, where the coefficients of genes can be considered together, we can also choose to correct the confounding factors first. The correction is mostly built upon the theory of how linear mixed model helps to correct confounding factors [81], and achieved by the rotation of $\mathbf{X}$ as following:
@@ Appendix G Challenges faced by existing methods on our benchmark
We tested these methods under three different settings: (i) providing general task instructions, (ii) providing detailed task instructions used by GenoAgent, and (iii) providing detailed task instructions and all necessary library functions as in GenoAgent. Each setting was tested on a subset of 50 GTA analysis problems. Our results show that none of the methods generated runnable code for preprocessing datasets downloaded from GEO. Persistent errors in the generated code prevented testable outputs, regardless of the level of detail provided.
@@ Appendix H Discussion on the limitations of GenoAgent
This section discusses the observed limitations of our baseline method, GenoAgent, on the GenoTEX benchmark. We identified that certain steps are inherently challenging, and instability in the feedback mechanism may hinder the agents’ iterative improvement process. Figures 4 and 4 illustrate the two types of feedback mechanisms in GenoAgent.